# Imaging the Freeze-out Source and Extracting Strong Interaction Parameters in Relativistic Heavy Ion Collisions

### v2

@@ IV Application to data
The direct imaging result in Fig. 4 (b) has important implications. First, Within the uncertainty, protons and antiprotons share the identical source distribution at the instant of freeze-out. In addition to the momentum distributions done in [29], the direct imaging result here completes the last piece of evidence of the matter-antimatter symmetry in coordinate space as they are produced and evolve in high energy HICs. This evidence is obtained without imposing assumptions to the shape of the source. Second, the profile of the reconstructed SFs deviates from the Gaussian distribution for both protons and antiprotons. To test whether the imaged source functions for $pp$ and $\bar{p}\bar{p}$ from STAR data are Gaussian distributions, each was fitted with a Gaussian function, and the corresponding chi-square (goodness of fit) values were obtained. For $pp$ and $\bar{p}\bar{p}$, the chi-square values were 91243.5 and 101.078, respectively, with 16 degrees of freedom (19 data points minus 3 parameters in the Gaussian fit). According to the chi-square test, quantitative calculation yields the probabilities for the SFs to be Gaussian is 00 and $2.2\times 10^{-14}$ for $p$ and $\bar{p}$, respectively. Thus, the Gaussian distribution hypothesis can be excluded. For a clear comparison, the Gaussian distribution obtained in [29] are shown by the curves in panel (b), respectively. It is evident that the emission source are more populated at smaller radius ($r<5$ fm) but less distributed at larger radius ($85<r<$ 8 fm), as zoomed in in the inset. It is consistent with the picture of an extremely fast collision. Namely, protons and antiprotons, formed by the coalescence of quarks and antiquarks created inside the overlapping region, are not fully randomized in coordinate space prior to the freeze-out instant, otherwise they are expected to extend to the outer space in Gaussian distribution. It offers a support to the recent idea of probing the initial shape of the nucleus via relativistic HIC [36, 37, 38, 39, 13, 40], since the collision is so rapid that the spatial information of the colliding nuclei is partially preserved.
@@
The impact of the bin-by-bin variation of the SF on the CF has been analyzed. It is observed that variations in the first few bins of the source function (except for the first bin which is below the low limit cut in LL model calculation) significantly impact the CF, whereas the variations at larger $r$ values are less influential. Inversely, the experimental uncertainty of the CFs near the peak ($q\approx 40~{}\rm MeV$) brings considerable variation to the SF. In order to achieve more accurate imaging, it is demanded to further reduce the uncertainty of the CF, particularly at small $q$. If given high-accuracy experimental CFs, the technique developed here offers a novel opportunity to identify exotic nuclear structure on future radioactive ion beam facilities. For instance, one can probe the proton density profile of bubble nucleus which has been recently proposed [41] but yet in debate [42]. Moreover, the imaging method developed in this work is compatible with machine learning techniques, offering a promising avenue for further enhancement and application in nuclear physics research.
V Summary
@@
A direct imaging method on femtometer scale has been realized by combining the Richardson-Lucy deblurring algorithm and the intensity interferometry in ultra relativistic heavy ion collisions. This method allows for a more accurate extraction of source distributions and final state interaction between particle pair based on the Lednický-Lyuboshitz model analysis. Applying this imaging method to $pp$ and $\bar{p}\bar{p}$ correlation functions in Au+Au at center-of-mass energy 200 GeV measured by the STAR collaboration, we observe the same freeze-out source functions for $p$ and $\bar{p}$ emissions with distinct non-Gaussian characteristics.These findings confirm the matter-antimatter symmetry in coordinate space once produced and suggest the ultrafast feature of the evolution of the collision, which is hardly accessible through other experimental means alone. The new method can be extended to three dimensional imaging and is inherently compatible with machine learning approaches. It represents a significant advancement of our ability to image the femtoscopic heavy ion collisions and to explore matter-antimatter physics in terrestrial laboratories.

### v3

@@ Abstract
The direct imaging result in Fig. 3 (b) has important implications. First, Within the uncertainty, protons and antiprotons share the identical source distribution at the instant of freeze-out. In addition to the momentum distribution analysis conducted in Ref. [47], the direct imaging results presented here provide the final piece of evidence supporting matter-antimatter symmetry in coordinate space, as it pertains to their production and evolution in high-energy HICs. Notably, this evidence is obtained without imposing any assumptions regarding the shape of the source. Second, the profile of the reconstructed SFs deviates from the Gaussian distribution for both protons and antiprotons. Quantitative calculation yields the probability for the SF to be Gaussian is 00 and $2.2\times 10^{-14}$ for $p$ and $\bar{p}$, respectively. For a clear comparison, the Gaussian distribution obtained in [47] are shown by the curves in panel (b), respectively. The emission source is more concentrated at smaller radii ($r<5$ fm) while showing reduced density at larger radii ($85<r<$ 8 fm), as highlighted in the inset. It is consistent with the picture of an extremely fast collision. Namely, protons and antiprotons, formed by the coalescence of quarks and antiquarks created inside the overlapping region, are not fully randomized in coordinate space prior to the freeze-out instant, otherwise they are expected to extend to the outer space in Gaussian distribution as a deduction of the central limit theorem. Our work offers a support to the recent idea of probing the initial shape and fluctuation of the nucleus via relativistic HIC [21, 54, 55, 56, 57, 58, 59, 60, 61], since the collision is so rapid that the spatial information of the colliding nuclei is partially preserved.
@@
The impact of the bin-by-bin variation of the SF on the CF has been analyzed. It is observed that variations in the first few bins of the source function (except for the first bin which is below the low limit cut in LL model calculation) significantly impact the CF, whereas the variations at larger $r$ values are less influential. Inversely, the experimental uncertainty of the CFs near the peak ($q\approx 40~{}\rm MeV$) brings considerable variation to the SF. In order to achieve more accurate imaging, it is demanded to further reduce the uncertainty of the CF, particularly at small $q$.
The method developed in this work holds significant potential for important applications. Given high-precision experimental CFs, this technique provides a novel approach to identify exotic nuclear structures, particularly in future radioactive ion beam facilities. For instance, it could be employed to probe the proton density profile of bubble nuclei—a recently proposed phenomenon [62, 63] that remains under debate[64]. Furthermore, when extended to three dimensions, this method could be applied at facilities such as RHIC to investigate higher-order deformations of atomic nuclei, which are typically studied through nuclear spectroscopy [65, 66].
@@
Summary - A direct imaging method on femtometer scale has been realized by combining the Richardson-Lucy deblurring algorithm and the intensity interferometry in relativistic heavy ion collisions. This method allows for a more accurate extraction of source distributions and final state interaction between particle pair based on the Lednický-Lyuboshitz model analysis. Applying this imaging method to $pp$ and $\bar{p}\bar{p}$ correlation functions in Au+Au at $\sqrt{S_{\rm NN}}=200$ GeV measured by the STAR collaboration, we observe the same freeze-out source functions for $p$ and $\bar{p}$ emissions with distinct non-Gaussian characteristics. These findings confirm the matter-antimatter symmetry in coordinate space once produced and suggest the ultrafast feature of the evolution of the collision, which is hardly accessible through other experimental means alone. This method can be extended to three-dimensional imaging for further applications and is inherently compatible with machine-learning approaches [67], opening new avenues for high-precision source reconstruction in heavy-ion collisions. It represents a significant advancement of our ability to image the femtoscopic heavy ion collisions and to explore matter-antimatter physics in terrestrial laboratories.